(a If the fifth-order maximum for pure-wavelength light falling on a double slit is at an angle of 41.8°, at what angle (in degrees) is the sixth-order maximum?

________ °

(b) What is the angle (in degrees) of the fourth minimum?

________ °

(c) What is the highest-order maximum possible here?
Answer #1

solution: for maxima in double slit experiment C dSinona no 1=1,2,... for minima in double slit experiment E 1 dsindr = (n+1)n=3 for fourth minima dsin cha = (n ++ ) d Sino = Sine - 82 SinX0-1834) Os sint 70.4689) 0 = sir To= 27.8340] e for making